A monopolist sells in two markets that have demand functions given by D1 (p1) = 100 - p1 and D2 (p2) = 100 - (1/2) p2: The marginal cost of production is constant at c = 20.

(a) Assume the firm charges different prices to each group. What will be the equilibrium quantities in markets 1 and 2?

(b) What market pays a higher price? Why?
Answer #1

a- a mmonopolist charges in such a fashion such that MR = MC

where MR = D(TR)/DQ, where TR = total revenue MR = marginal revenue

so in market 1

100-2Q = 20 AND Q = 40, THIS P1 = 60

IN MARKET 2

MR2 = MC

200-4Q2 = 20

Q2 = 45, THUS P2 = 110

B- SO MARKET 2 PAYS HIGHER PRICES BECAUSE MARKET 2 IS MORE INELASTIC DEMANDED , THUS CHARGING HIGHER PRICES THERE WONT DECREASE THE QUANTITY MUCH.